OpenStudy (anonymous):

Explain, in complete sentences, how adding and subtracting radicals and complex numbers is similar to adding and subtracting polynomials. Give one example of an addition or subtraction problem of radicals and one example of an addition or subtraction problem of complex numbers.

OpenStudy (anonymous):

it is similar in that you are combining like terms. for example \[2+\sqrt{5}+4-3\sqrt{5}\] combine as \[6-2\sqrt{5}\] just as \[2+x+4-3x\] combine to give \[6-2x\]